Find the solution of the initial value problem y′′+4y=t^2+6e^t, y(0)=0, y′(0)=5.
Enter an exact answer.
Enclose arguments of functions in parentheses. For example, sin(2x).
1
Expert's answer
2020-07-09T19:32:39-0400

The given equation can be written as, "\left(D^{2}+4\right) y=t^{2}+6e^{t}". The auxiliary equation is, "m^{2}+4=0". Solving, we get "m=\pm 2i". The complementary function is,

"C.F=c_{1} \cos (2 t)+c_{2} \sin (2 t)." The particular integral is given by,

"\begin{aligned}
P.I &=\frac{1}{D^{2}+4}\left(t^{2}+6 e^{t}\right) \\
&=\frac{1}{D^{2}+4} \cdot t^{2}+\frac{1}{D^{2}+4} \cdot 6 e^{t} \\
&=\frac{1}{4}\left(1+\frac{D^{2}}{4}\right)^{-1} \cdot t^{2}+\frac{6 e^{t}}{5} \\
&=\frac{1}{4}\left(1-\frac{D^{2}}{4}+\cdots\right) \cdot t^{2}+\frac{6 e^{t}}{5} \\
&=\frac{1}{4}\left(t^{2}-\frac{1}{2}\right)+\frac{6 e^{t}}{5} \\
&=\frac{t^{2}}{4}+\frac{6 e^{t}}{5}-\frac{1}{8}
\end{aligned}"

The general solution is, "\displaystyle y=c_{1} \cos (2 t)+c_{2} \sin (2 t)+\frac{t^{2}}{4}+\frac{6 e^{t}}{5}-\frac{1}{8}".

Hence, "\displaystyle y'=-2 c_{1} \sin (2 t)+2 c_{2} \cos (2 t)+\frac{t}{2}+\frac{6 e^{t}}{5}" .


Given the initial conditions, "y(0)=0, y^{\prime}(0)=5". Using the initial conditions we get,


"\begin{aligned}
\displaystyle y(0) & = c_{1} +\frac{6 }{5}-\frac{1}{8}\\
0&=c_{1} +\frac{6 }{5}-\frac{1}{8}\\
c_{1}&=-\frac{43}{40}
\end{aligned}"

"\begin{aligned}
\displaystyle y'(0) & = 2c_{2} +\frac{6 }{5}\\
5&=2c_{2} +\frac{6}{5}\\
2c_{2}&= 5-\frac{6}{5}\\
c_{2}&=\frac{19}{10}
\end{aligned}"

Hence, the general solution is,

"\begin{aligned}
y&=\frac{19}{10} \sin (2 t)-\frac{43}{40} \cos (2 t)+\frac{t^{2}}{4}+\frac{6 e^{t}}{5}-\frac{1}{8}\\
y&=\frac{1}{40}\bigg(76 \sin (2 t)-43 \cos (2 t)+10 t^{2}+48 e^{t}-5\bigg)
\end{aligned}"